OpenStudy (anonymous):

Hi guys, I am not using you as a homework help. My mom put me in a computer class and this question is really confusing, please tell me how to do it. How do you change powers with negative exponents to powers with positive exponents?

OpenStudy (anonymous):

can u give an example with negative exponets?

OpenStudy (anonymous):

3^-2

OpenStudy (anonymous):

a^(-n) = 1/a^n for a and n non zero

OpenStudy (anonymous):

@xxferrocixx is right...

OpenStudy (anonymous):

@xxferrocixx how?

OpenStudy (anonymous):

I am confused, how is this related to my question?

OpenStudy (anonymous):

are you guys still here? sorry I went away, my internet crashed.

OpenStudy (anonymous):

because that is the law for negative exponent. so 3^-2 = 1/(3^2) = 1/9

OpenStudy (anonymous):

Ok, I kind of get it now, I am still confused on why you do 1/(3^2) to make it positive

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Ok, thank you for your help guys. I have the basic concept now, will try to wing it.